Factory Overhead Cost Variances

Thomas Textiles Corporation began November with a budget for 42,000 hours of production in the Weaving Department. The department has a full capacity of 56,000 hours under normal business conditions. The budgeted overhead at the planned volumes at the beginning of November was as follows:

Variable overhead $126,000
Fixed overhead 89,600
Total $215,600

The actual factory overhead was $218,200 for November. The actual fixed factory overhead was as budgeted. During November, the Weaving Department had standard hours at actual production volume of 44,000 hours.

Determine the variable factory overhead controllable variance and the fixed factory overhead volume variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number. Round your interim computations to the nearest cent, if required.

a. Variable factory overhead controllable variance: $  Favorable

b. Fixed factory overhead volume variance: $ Unfavorable
Answer #1

Solution:

variable overhead Fixed overhead
Budgeted Total 126000 89600
/Planned hours 42000 56000
Standard rates 3.00 1.60

Actual cost of variable overhead = Total - Fixed overehad = 218200 - 89600 = $128,600

Variable factory overhead controllable variance = Actual cost of variable overhead - Standard cost of variable overhead

= $128600 - (44000*$3.00) = - $3,400 F

Fixed factory overhead volume variance = Budgeted fixed overhead - Fixed overhead applied

= $89600 - (44000 * $1.60) = $19,200 U